Juice12231:

Bart opened a credit card at a department store with a APR of 19.99% compounded monthly.what is the APY on this credit card?

Juice12231:

A.39.98% B.26.81% C.21.93% D.1.67%

jimthompson5910:

The formula you use is \(\large \text{APY} = \left(1 + \frac{\text{APR}}{n}\right)^{n}-1\)


Here's a similar example: You apply for a credit card that has 24% APR and it's compounded monthly. What is the APY? APR = 24% = 0.24 We compound monthly, so the money is compounded n = 12 times a year \(\large \text{APY} = \left(1 + \frac{\text{APR}}{n}\right)^{n}-1\) \(\large \text{APY} = \left(1 + \frac{\text{0.24}}{12}\right)^{12}-1\) Note how I'm using the decimal form of 24% \(\large \text{APY} = \left(1 + 0.02\right)^{12}-1\) \(\large \text{APY} = \left(1.02\right)^{12}-1\) \(\large \text{APY} \approx 1.26824179 -1\) \(\large \text{APY} \approx 0.26824179\) \(\large \text{APY} \approx 0.2682\) \(\large \text{APY} \approx 26.82\%\) This isn't the answer to your question as it's just an example of course. Use these steps to find the APY when APR = 19.99%